A building has a rectangular room and a square room with equal floor space. The rectangular room is 3 meters longer and 2 meters narrower than the square room. Find the floor space of each. Please explain the steps in finding the answer as well.
1
Expert's answer
2011-06-17T07:27:01-0400
Let's denote the side of the square room as x. Then the one side of rectangular room is x +3, the other side is x - 2.
Sinse both rooms have an equal area,
x*x = (x+3)(x-2)
x2 = x2 + x - 6
x = 6.
Thus the floor space of each room is 62 = 36 m2.